# On the nullspace of split graphs

### v1

@@ 3 Null space of split graphs
    The condition $\deg_{Sp}(x)=\deg_{Sp}(y)$ for all $x,y\in S$ means that all vertices in $S$ have the same number of neighbors in $K$ (since vertices in $S$ have no neighbors in $S$). Thus, there exists $d\geq 0$ such that $|N_{Sp}(x)\cap K|=d$ for all $x\in S$.
    By connectivity and the fact that $K$ is nonempty, we have $d\geq 1$. Therefore, each column sum of $R$ equals $d$, which means each row sum of $R^{t}$ equals $d$, so:
@@
    which shows that $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R^{t})$. By Lemma 3.1 with $A=R$ and $B=J=\operatorname{1\!\!1}\operatorname{1\!\!1}^{t}$, this implies $\operatorname{\mathcal{N}}(R)\subset\operatorname{\mathcal{N}}(J)$. ∎
@@
The condition $K=\dot{\bigcup}_{v\in S}N_{Sp}(v)$ means that the neighborhood sets $\{N_{Sp}(v)\cap K:v\in S\}$ form a partition of $K$. This implies that each vertex in $K$ has exactly one neighbor in $S$, so every row of $R$ contains exactly one nonzero entry (which is 1).
Consequently, all vertices in $K$ have the same degree in $Sp$, since each has degree |K|1|K|-1 within $K$ plus exactly 1 neighbor in $S$, giving $\deg_{Sp}(u)=|K|$ for all $u\in K$. By Theorem 3.6(1), this implies $\operatorname{Supp}(Sp)\subset S$. Then by Theorem 3.4(3), we have $\operatorname{Cker}{\left(Sp\right)}=\{0\}$, and by Theorem 3.2(3), $\operatorname{nl}(Sp)=\operatorname{nl}(R)$.
@@ 4 Unbalanced split graphs
From now on, it is useful to keep in mind this straightforward observation: if $G$ is a graph and if $K,S$ are, respectively, a clique and an independent set of $G$, then $|K\cap S|\in\{0,1\}$. Moreover, we strongly recommend to see [nordhausgaddumcheng2016split] and [whitman2020split] (page 8) before reading this section.
@@
Let $Sp(K,S)$ be a split graph of order $n$. If $\omega=\omega(Sp)$ and $\alpha=\alpha(Sp)$, then we have the following:
@@
    $\omega+\alpha\in\{n,n+1\}$;
@@
    $|K|\geq\omega-1$ and $|S|\geq\alpha-1$;
  3. 3.
    $(|K|,|S|)\in\{(\omega,\alpha),(\omega,\alpha-1),(\omega-1,\alpha)\}$;
  4. 4.
    $(|K|,|S|)\in\{(\omega,\alpha-1),(\omega-1,\alpha)\}$ if and only if there exist a maximum clique $K_{\max}$, a maximum independent set $S_{\max}$ and a vertex $w$ in $Sp$ such that
    $V(Sp)=K_{\max}\cup S_{\max},\ K_{\max}\cap S_{\max}=\{w\},\ N_{Sp}(w)=K-w.$
  5. 5.
    there exist a maximum clique $K_{\max}$ and a maximum independent set $S_{\max}$ in $Sp$ such that $V(Sp)=K_{\max}\cup S_{\max}$;
  6. 6.
    $(|K|,|S|)\!=\!(\omega,\alpha)$ if and only if $(K,S)$ is the unique s-partition for $Sp$.
@@
  1. (1)
    Let $K_{\max}$ and $S_{\max}$ be respectively a maximum clique and a maximum independent set of $Sp$. Since $V(Sp)\supseteq K_{\max}\cup S_{\max}$, we have
    $n+|K_{\max}\cap S_{\max}|\geq|K_{\max}|+|S_{\max}|=\omega+\alpha.$
    Since $|K_{\max}\cap S_{\max}|\in\{0,1\}$, we infer that $n+1\geq\omega+\alpha$. On the other hand, recall that $\omega\geq|K|$ and $\alpha\geq|S|$. Then,
    $\omega+\alpha\geq|K|+|S|=n.$
  2. (2)
    If $|K|\leq\omega-2$, then by using (1) we obtain
    $\omega+\alpha-1\leq|K|+|S|\leq\omega-2+|S|,$
    and so $|S|\geq\alpha+1$, which is impossible. Hence, $|K|\geq\omega-1$. Similar arguments can be applied if $|S|\leq\alpha-2$.
  3. (3)
    If $(|K|,|S|)=(\omega-1,\alpha-1)$, then $n=|K|+|S|=\omega+\alpha-2$, which contradicts (1). Now, the result follows immediately from (2).
  4. (4)
    If $(|K|,|S|)=(\omega-1,\alpha)$, recall that every clique in $Sp$ intersects $S$ in at most one vertex. Therefore, a clique of size $|K|+1$ must be of the form $K\dot{\cup}w$, for some $w\in S$ such that $N_{Sp}(w)=K=K-w$. Consequently, we define $K_{\max}=K\dot{\cup}w$ and $S_{\max}=S$.
    If $(|K|,|S|)=(\omega,\alpha-1)$, then an independent set of size $|S|+1$ must be of the form $S\dot{\cup}w$, for some $w\in K$ such that $N_{Sp}(w)=K-w$. We define $K_{\max}=K$ and $S_{\max}=S\dot{\cup}w$.
    For the converse, assume that we have $K_{\max},S_{\max}$ and $w$ with those properties, but there is a bipartition $(K^{\prime},S^{\prime})$ for $Sp$ such that
    $(|K^{\prime}|,|S^{\prime}|)\notin\{(\omega,\alpha-1),(\omega-1,\alpha)\}.$
    Then, $(|K^{\prime}|,|S^{\prime}|)=(\omega,\alpha)$, by (3). Since $(K_{\max}-w,S_{\max})$ is also a bipartition for $Sp$, we have
    $\omega+\alpha=|K^{\prime}|+|S^{\prime}|=n=|K_{\max}-w|+|S_{\max}|=\omega-1+\alpha,$
    which is absurd. Therefore, every bipartition $(K,S)$ for $Sp$ has to satisfy that $(|K|,|S|)\in\{(\omega,\alpha-1),(\omega-1,\alpha)\}$.
  5. (5)
    From (1) we know that $\omega+\alpha\in\{n,n+1\}$. If $\omega+\alpha=n=|K|+|S|$, it immediately follows from (3) that $(|K|,|S|)=(\omega,\alpha)$. Hence, define $K_{\max}=K$ and $S_{\max}=S$. If $\omega+\alpha=n+1=|K|+|S|+1$, then
    $(|K|,|S|)\in\{(\omega-1,\alpha),(\omega,\alpha-1)\},$
    by (3). Now, use (4).
  6. (6)
    Suppose that $(K,S)$ is the unique bipartition for $Sp$, and use item (5) to write $V(Sp)=K_{\max}\cup S_{\max}$. If $K_{\max}\cap S_{\max}=\{w\}$, we see that
    $(K_{\max},S_{\max}-w),\ (K_{\max}-w,S_{\max})$
    are bipartitions for $Sp$. However, by uniqueness of $(K,S)$ we deduce that
    $K_{\max}-w=K=K_{\max},$
    which is absurd. Therefore, $K_{\max}\cap S_{\max}=\varnothing$, which means that $(K_{\max},S_{\max})$ is a bipartition for $Sp$. Since $(K,S)$ is unique, it must be $K_{\max}=K$ and $S_{\max}=S$. Hence, $(|K|,|S|)=(\omega,\alpha)$.
    Conversely, assume that $(|K|,|S|)=(\omega,\alpha)$. If $(K^{\prime},S^{\prime})$ is any other bipartition for $Sp$, then
    $\omega+\alpha=|K|+|S|=n=|K^{\prime}|+|S^{\prime}|,$
    and so $(|K^{\prime}|,|S^{\prime}|)=(\omega,\alpha)$ as well, by (3). Consequently, $(K^{\prime},S^{\prime})\neq(K,S)$ implies $|K-K^{\prime}|=|K^{\prime}-K|\neq 0$. Let $x\in K-K^{\prime}$. Since $x\in K$, it is easy to see that $\deg_{Sp}(x)\geq\omega$ (otherwise, $S\dot{\cup}x$ would be an independent set of size $\alpha+1$). On the other hand, since $x\in S^{\prime}$, we must have $\deg_{Sp}(x)\leq\omega-1$ (otherwise, $K^{\prime}\dot{\cup}x$ would be a clique of size $\omega+1$). Hence, $\omega\leq\deg_{Sp}(x)\leq\omega-1$, which is absurd. Therefore, $(K^{\prime},S^{\prime})=(K,S)$. ∎
A split graph $Sp(K,S)$ is said to be balanced if $|K|=\omega(Sp)$ and $|S|=\alpha(Sp)$. Otherwise, we say that $Sp$ is unbalanced (or not balanced). Then, by item (6) of Proposition 4.1 we see that $Sp$ is balanced if and only if it has a unique bipartition. As a consequence, unbalanced split graphs are exactly those with two or more s-partitions. From items (4) and (6) of Proposition 4.1 we can also deduce that $Sp$ is unbalanced if and only if
$(|K|,|S|)\in\{(\omega,\alpha-1),(\omega-1,\alpha)\}.$
We say that a vertex $w$ of a split graph $Sp$ is swing if there exist s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$ and $w\in S^{\prime}$. Let $W(Sp)$ be the set of all swing vertices of $Sp$. Then, it follows that $Sp$ is balanced if and only if $W(Sp)=\varnothing$. However, SpW(Sp)Sp-W(Sp) may not be balanced in general. From item (4) of Proposition 4.1 we deduce that $w\in W(Sp)$ if and only if there are s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$, $w\in S^{\prime}$ and
$K-w=N_{Sp}(w)=K^{\prime},\ (|K|,|S|)=(\omega,\alpha-1),\ (|K^{\prime}|,|S^{\prime}|)=(\omega-1,\alpha).$
If $S\!P(Sp)=\{(K,S)\}$ is the collection of all the s-partitions for $Sp$, then
$V(Sp)=W(Sp)\ \dot{\cup}\ \bigcap_{(K,S)\in S\!P(Sp)}K\ \dot{\cup}\ \bigcap_{(K,S)\in S\!P(Sp)}S.$
@@
The following theorem states that in unbalanced split graphs where the independent set is maximum-sized, all support vertices must lie in the independent set. This has significant implications for understanding the nullspace structure of unbalanced split graphs.
@@
If $Sp(K,S)$ is an unbalanced split graph with $|S|=\alpha(Sp)$, then $\operatorname{Supp}(S)\subset S$.
@@
Since $Sp$ is unbalanced with $|S|=\alpha(Sp)$, by Proposition 4.1(3) and (4), we must have $(|K|,|S|)=(\omega-1,\alpha)$. Then by Proposition 4.1(4), there exists a vertex $w\in S$ such that $N_{Sp}(w)=K$.
In the bipartite adjacency matrix $R$, this means that the column corresponding to $w$ (i.e., $R_{*w}$) has all entries equal to 1, so $R_{*w}=\operatorname{1\!\!1}$. Therefore, $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R)$, and by Theorem 3.2(5), we conclude that $\operatorname{Supp}(Sp)\subset S$. ∎
A threshold graph is a graph that can be constructed from $K_{1}$ by repeatedly applying either of the following two operations: 1) addition of an isolated vertex; 2) addition of a universal vertex. Thus, each unlabeled threshold graph of order $n$ can be uniquely identified with a binary sequence $0b_{2}b_{3}\ldots b_{n}$, with $b_{i}\in\{0,1\}$, where 0=0=“add an isolated vertex” and 1=1=“add a universal vertex”. Threshold graphs are characterized as those graphs that do not contain induced subgraphs isomorphic to $P_{4}$, $C_{4}$ or $2K_{2}$ ([barrus.west.A4], page 2). Since a graph is split if and only if it does not contain induced subgraphs isomorphic to $2K_{2}$, $C_{4}$ or $C_{5}$ ([barrus.west.A4], page 14), it is clear that threshold graphs are split. If $0b_{2}\ldots b_{n}$ is the binary sequence associated with a threshold graph $Sp$, it is easy to see that
@@
is a bipartition for $Sp$. Then, $N_{Sp}(v_{1})=K$, which shows that $Sp$ is unbalanced. The following corollary is known and it can be found in [panelo5316214core].
@@
If $Sp(K,S)$ is threshold and $|S|=\alpha(Sp)$, then $\operatorname{Supp}(Sp)\subset S$.
@@
Let $Sp$ be a split graph and let $W(Sp)$ be the set of swing vertices of $Sp$. If $w\in W(Sp)$, then $[w]=W(Sp)$.
@@
If $W(Sp)=\{w\}$, the inclusion $W(Sp)\subset[w]$ is obvious. If $|W(Sp)|\geq 2$, let $w^{\prime}\in W(Sp)-\{w\}$. Since $w$ and $w^{\prime}$ are swing, there are bipartitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that
$w\in S,\ w^{\prime}\in S^{\prime},\ |S|=\alpha(Sp)=|S^{\prime}|,\ N_{Sp}(w)=K,\ N_{Sp}(w^{\prime})=K^{\prime}.$
Suppose that $K\neq K^{\prime}$. If $x\in K-K^{\prime}$, then $x\in S^{\prime}$. Hence, $x\in W(Sp)$, and so $K^{\prime}=N_{Sp}(x)=K-x$. Now, observe that $S^{\prime}=S\dot{\cup}\{x\}$: this is impossible because $S$ and $S^{\prime}$ are maximum independent sets. Therefore, $K=K^{\prime}$, which means that $w^{\prime}$ and $w$ are twins. We have just proven the inclusion $W(Sp)\subset[w]$.
To prove the reverse inclusion, consider a vertex $w^{\prime}$ with $N_{Sp}(w^{\prime})-w=N_{Sp}(w)-w^{\prime}$. Since $w$ is swing, there is a bipartition $(K,S)$ for $Sp$ such that $w\in S$ and $N_{Sp}(w)=K$. Then, $N_{Sp}(w^{\prime})=K-w^{\prime}$, which means that $w^{\prime}\in W(Sp)$. Hence, $w]\subset W(Sp)$. ∎
@@
Since $x\in\ker(G)$, we have
$G_{v*}x=x_{u}+\sum_{i\in N_{v}-u}x_{i}=0=G_{u*}x=x_{v}+\sum_{i\in N_{u}-v}x_{i}.$
Since N(v)u=N(u)vN(v)-u=N(u)-v, we have
$\sum_{i\in N(v)-u}x_{i}=\sum_{i\in N(u)-v}x_{i}.$
@@
Let $Sp(K,S)$ be an unbalanced split graph and let $W(Sp)$ be the set of its swing vertices. Then, we have the following:
@@
    $(|K|,|S|)\in\{(\omega(Sp),\alpha(Sp)-1),(\omega(Sp)-1,\alpha(Sp))\}$;
@@
    $W(Sp)$ is a clique or an independent set of mutually twin vertices;
@@
    if $W(Sp)=\{w\}$, then $S\cup\{w\}$ is the unique maximum independent set of $Sp$ and $\operatorname{Supp}(Sp)\subset S\cup\{w\}$;
@@
    if $W(Sp)$ is a clique of size $\geq 2$, then
    $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing,\ \operatorname{Supp}(Sp)\subset S\setminus W(Sp);$
  5. 5.
@@
    $W(Sp)\subset\operatorname{Supp}(Sp)\subset S\cup W(Sp).$
@@
  1. (1)
    We already know from previous discussions that this is equivalent to being unbalanced.
@@
    By using item (5) of Proposition 4.1, we can write $V(Sp)=K_{\max}\cup S_{\max}$. Since these two sets have to share a swing vertex and since $|W(Sp)|=1$, it follows that in $Sp$, $K_{\max}$ is the unique maximum clique and $S_{\max}$ is the unique maximum independent set. Hence, $Sp$ admits exactly two s-partitions: $(K_{\max},S_{\max}\setminus\{w\})$, $(K_{\max}\setminus\{w\},S_{\max})$. From this, it becomes clear that $S_{\max}=S\cup\{w\}$. Taking $(K,S)=(K_{\max}\setminus\{w\},S_{\max})$, we are now able to apply Theorem 4.2.

### v2

@@ 3 Nullspace of split graphs
    The condition $\deg(x)=\deg(y)$ for all $x,y\in S$ means that all vertices in $S$ have the same number of neighbors in $K$ (since vertices in $S$ have no neighbors in $S$). Thus, there exists $d\geq 0$ such that $|N(x)\cap K|=d$ for all $x\in S$.
    By connectedness and the fact that $K$ is nonempty, we have $d\geq 1$. Therefore, each column sum of $R$ equals $d$, which means each row sum of $R^{t}$ equals $d$, so:
@@
    which shows that $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R^{t})$. By Lemma 3.1, this implies $\operatorname{\mathcal{N}}(R)\subset\operatorname{\mathcal{N}}(J)$. ∎
@@
The condition $K=\dot{\bigcup}_{v\in S}N(v)$ means that the neighborhood sets $\{N(v):v\in S\}$ form a partition of $K$. This implies that each vertex in $K$ has exactly one neighbor in $S$, so every row of $R$ contains exactly one nonzero entry (which is 1). Consequently, all vertices in $K$ have the same degree in $Sp$, since each has degree |K|1|K|-1 within $K$ plus exactly 1 neighbor in $S$, giving $\deg(u)=|K|$ for all $u\in K$. By Theorem 3.6(1), this implies $\operatorname{Supp}(Sp)\subset S$. Then by Theorem 3.4(3), we have $\operatorname{Cker}{\left(Sp\right)}=\{0\}$, and by Theorem 3.2(3), $\operatorname{nl}(Sp)=\operatorname{nl}(R)$.
@@ 4 Unbalanced split graphs
A split graph $Sp(K,S)$ is said to be balanced if $|K|=\omega(Sp)$ and $|S|=\alpha(Sp)$. Otherwise, we say that $Sp$ is unbalanced (or not balanced). It is known that (see [3], [9] and [8]):
  1. 1.
    SpSp is balanced if and only if it has a unique s-partition (thus, unbalanced split graphs are exactly those with two or more s-partitions);
  2. 2.
    SpSp is balanced if and only if $\omega(Sp)+\alpha(Sp)=|Sp|$.
  3. 3.
    $Sp(K,S)$ is unbalanced if and only if
    $(|K|,|S|)\in\{(\omega(Sp),\alpha(Sp)-1),(\omega(Sp)-1,\alpha(Sp))\}.$
  4. 4.
    SpSp is unbalanced if and only if $\omega(Sp)+\alpha(Sp)=|Sp|+1$.
We say that a vertex $w$ of a split graph $Sp$ is swing if there exist s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$ and $w\in S^{\prime}$. Let $W(Sp)$ be the set of all swing vertices of $Sp$. Then, it follows that $Sp$ is balanced if and only if $W(Sp)=\varnothing$. However, $Sp\setminus W(Sp)$ may not be balanced in general. It is easy to see that $w\in W(Sp)$ if and only if there are s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that $w\in K$, $w\in S^{\prime}$ and
$K\setminus\{w\}=N_{Sp}(w)=K^{\prime},$ (1)
$(|K|,|S|)=(\omega(Sp),\alpha(Sp)-1),\ (|K^{\prime}|,|S^{\prime}|)=(\omega(Sp)-1,\alpha(Sp)).$
If $Sp$ is a split graph, let:
  1. 1.
    $\mathcal{K}(Sp)$ be the set of all cliques $K$ of $Sp$ such that $(K,S)$ is an s-partition for $Sp$, for some $S$;
  2. 2.
    $\mathcal{S}(Sp)$ be the set of all independent sets $S$ of $Sp$ such that $(K,S)$ is an s-partition for $Sp$, for some $K$;
  3. 3.
    $K^{*}(Sp)$ be the set of vertices that are “always clique” in $Sp$, i.e., $K^{*}(Sp)=\bigcap(K\in\mathcal{K}(Sp))$;
  4. 4.
    $S^{*}(Sp)$ be the set of vertices that are “always independent” in $Sp$, i.e., $S^{*}(Sp)=\bigcap(S\in\mathcal{S}(Sp))$.
Then, by the definition of swing vertex, it is clear that
$V(Sp)=K^{*}(Sp)\ \dot{\cup}\ W(Sp)\ \dot{\cup}\ S^{*}(Sp).$ (2)
The concept of swing vertex lies implicitly among several works on split graphs. But, as far as we know, the term “swing” first appears explicitly in [21] (page 8). However, there the definition of swing vertex is slightly different: a vertex $w$ is swing in $Sp(K,S)$ if
$N_{Sp}(w)=K\setminus\{w\}.$ (3)
Although (3) agrees with (1), we think there is an issue with (3): it depends on the s-partition. To see that, consider the split graph $Sp=abc\approx P_{3}$ with the s-partitions $(K,S)=(\{a,b\},c)$ and $(K^{\prime},S^{\prime})=(a,\{b,c\})$. By our definition, $W(Sp)=\{a,c\}$, which is an invariant of $Sp$. However, using (3), we observe that: 1) in $Sp(K,S)$, $a$ is swing but $c$ is not; 2) in $Sp(K^{\prime},S^{\prime})$, $c$ is swing but $a$ is not. For this reason, we decided to modify (3). As we have seen above, this choice allows us to immediately obtain the useful vertex-decomposition (2).
Next theorem is a translation to our setting of a known result that characterizes the s-partitions of an unbalanced split graph.
Theorem 4.1 (Cheng, Collins, Trenk, [3], page 2348).
Let $Sp$ be an unbalanced split graph and let $\mathcal{KS}$ be the collection of all the s-partitions for $Sp$. Then, we have the following (for simplicity, we will use $W,K^{*},S^{*}$ instead of $W(Sp),K^{*}(Sp),S^{*}(Sp)$).
  1. 1.
    If $W=\{w\}$, then
    $\mathcal{KS}=\{(K^{*}\cup w,S^{*}),(K^{*},S^{*}\cup w)\};$
    moreover, $K^{*}\cup w$ is the unique maximum clique of $\mathcal{K}(Sp)$ and $S^{*}\cup w$ is the unique maximum independent set of $\mathcal{S}(Sp)$.
  2. 2.
    If $W$ is a clique of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*}\cup W,S^{*})\}\cup\{(K^{*}\cup(W\setminus\{w\}),S^{*}\cup w):w\in W\};$
    moreover, $K^{*}\cup W$ is the unique maximum clique of $\mathcal{K}(Sp)$ and $|S^{*}\cup w|=\alpha(Sp)$.
  3. 3.
    If $W$ is an independent set of size $\geq 2$, then
    $\mathcal{KS}=\{(K^{*},S^{*}\cup W)\}\cup\{(K^{*}\cup w,S^{*}\cup(W\setminus\{w\})):w\in W\};$
    moreover, $S^{*}\cup W$ is the unique maximum independent set of $\mathcal{S}(Sp)$ and $|K^{*}\cup w|=\omega(Sp)$.
The following proposition is a refinement of Theorem 4.1, because, essentially, tells us that $\mathcal{K}(Sp)$” and $\mathcal{S}(Sp)$” can be replaced with “SpSp” in Theorem 4.1.
Theorem 4.2.
Let $Sp$ be an unbalanced split graph. Then,
  1. 1.
    $\mathcal{K}(Sp)$ contains all the maximum cliques of $Sp$.
  2. 2.
    $\mathcal{S}(Sp)$ contains all the maximum independent sets of $Sp$.
Proof.
  1. (1).
    Let $K$ be a maximum clique of $Sp$. Since $Sp$ is unbalanced, we know that $\omega(Sp)+\alpha(Sp)=|Sp|+1$, and so $\alpha-1=|Sp|-\omega=|V(Sp)\setminus K|$. Let $S$ be a maximum independent set of $Sp$. Since $|K\cap S|\leq 1,|S|=\alpha$ and $|V(Sp)\setminus K|=\alpha-1$, it must be $S=(V(Sp)\setminus K)\dot{\cup}w$, for some $w\in K$. Consequently, $V(Sp)\setminus K$ is an independent set of $Sp$. Thus, $(K,V(Sp)\setminus K)$ is an s-partition for $Sp$, which means that $K\in\mathcal{K}(Sp)$.
  2. (2).
    Let $S$ be a maximum independent set of $Sp$. Then, $S$ is a maximum clique of the graph $\overline{Sp}$ (the complement of $Sp$), which is clearly split and unbalanced. From (1), we know that $S\in\mathcal{K}(\overline{Sp})$, which means that $V(\overline{Sp})\setminus S$ is an independent set of $\overline{Sp}$. Hence, $V(Sp)\setminus S$ is a clique of $Sp$, which shows that $S\in\mathcal{S}(Sp)$.
Surprisingly, Theorem 4.2 does not hold in general for balanced split graph. In fact, every balanced $Sp(K,S)$ with an independent vertex of degree |K|1|K|-1 is a counterexample ($P_{4}$ is the “simplest” one). In other words, a balanced split graph $Sp(K,S)$ has always a unique s-partition, but $K$ might not be the only maximum clique of $Sp$. An analogous phenomenon can be observed for maximum independent sets. The following result clarifies this behavior of the balanced split graphs.
@@
Let $Sp(K,S)$ be a balanced split graph. Then, we have the following.
@@
    $K$ is the only maximum clique of $Sp$ if and only if $Sp$ does not contain an independent vertex of degree |K|1|K|-1.
@@
    $S$ is the only maximum independent set of $Sp$ if and only if $Sp$ does not contain a clique vertex $x$ such that $|N_{Sp}(x)\cap S|=1$.
@@
  1. (1).
    If $x$ is a vertex of $Sp$ of degree |K|1|K|-1, then $x$ must be independent (otherwise, $x$ would be swing). Now, observe that $N(x)\cup x$ is a maximum clique of $Sp$, different from $K$.
    For the converse, suppose that $K^{\prime}$ is a maximum clique of $Sp$, different from $K$. Since $|K^{\prime}\cap S|\leq 1$, it must be $K^{\prime}=(K\setminus\{w\})\dot{\cup}\{x\}$, for some $w\in K$ and $x\in S$. Since $x\in K^{\prime}$, we have $K\setminus\{w\}\subset N(x)$. Since $\in S, xx$ is not adjacent to any vertex of $S$, which means that $N(x)\subset K$. If $w\in N(x)$, then $N(x)=K$ and hence $x\in W(Sp)=\varnothing$. This absurdity implies $N(x)=K\setminus\{w\}$, i.e., $\deg(x)=|K|-1$.
  2. (2).
    If $x$ is a vertex of $Sp$ such that $N(x)\cap S=\{w\}$ for some $w$, it is clear that $x\in K$ and $(I\setminus\{w\})\dot{\cup}\{x\}$ is a maximum independent set different from $S$.
    For the converse, suppose that $S^{\prime}$ is a maximum independent set of $Sp$, different from $S$. Since $|K\cap S^{\prime}|\leq 1$, it must be $S^{\prime}=(S\setminus\{w\})\dot{\cup}\{x\}$, for some $w\in S$ and $x\in K$. Since $x\in K$, $K\setminus x\subset N(x)$. Since $\in S^{\prime}, xx$ is not adjacent to any vertex of $S\setminus\{w\}$. Notice that $w\in N(x)$ (otherwise, $x$ would be swing). Therefore, $N(x)=(K\setminus x)\dot{\cup}\{w\}$, which means that $|N(x)\cap S|=1$.
@@
The following theorem states that in unbalanced split graphs, supported vertices must lie in the intersection of all the maximum independent sets. This has significant implications for understanding the nullspace structure of unbalanced split graphs.
@@
Let $Sp$ be an unbalanced split graph. Then, $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets of $Sp$.
@@
Consider an arbitrary s-partition $(K,S)$ for $Sp$ such that $|S|=\alpha(Sp)$. The existence of such is guaranteed by Theorem 4.1. Then, by Theorem 4.1, there exists a vertex $w\in S$ such that $N(w)=K$, which means that $R_{*w}=\operatorname{1\!\!1}$. Since $\operatorname{1\!\!1}\in\operatorname{\mathcal{R}}(R)$, we conclude by Theorem 3.2(5) that $\operatorname{Supp}(Sp)\subset S$. Therefore, $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets in $\mathcal{S}(Sp)$. Now, use Theorem 4.2(2). ∎
A threshold graph is a graph that can be constructed from $K_{1}$ by repeatedly applying either of the following two operations: 1) addition of an isolated vertex; 2) addition of a universal vertex. Thus, each unlabeled threshold graph of order $n$ can be uniquely identified with a binary sequence $0b_{2}b_{3}\ldots b_{n}$, with $b_{i}\in\{0,1\}$, where 0=0=“add an isolated vertex” and 1=1=“add a universal vertex”. Threshold graphs are characterized as those graphs that do not contain induced subgraphs isomorphic to $P_{4}$, $C_{4}$ or $2K_{2}$ ([2], page 2). Since a graph is split if and only if it does not contain induced subgraphs isomorphic to $2K_{2}$, $C_{4}$ or $C_{5}$ ([2], page 14), it is clear that threshold graphs are split. If $0b_{2}\ldots b_{n}$ is the binary sequence associated with a threshold graph $Sp$, it is easy to see that
@@
is an s-partition for $Sp$. Then, $N_{Sp}(v_{1})=K$, which shows that $Sp$ is unbalanced. The following corollary is a refined version of a result that can be found in [14].
@@
If $Sp$ is threshold, then $\operatorname{Supp}(Sp)$ is contained in the intersection of all the maximum independent sets of $Sp$.
@@
Let $Sp$ be a split graph. If $w\in W(Sp)$, then $[w]=W(Sp)$.
@@
If $W(Sp)=\{w\}$, the inclusion $W(Sp)\subset[w]$ is obvious. If $|W(Sp)|\geq 2$, let $w^{\prime}\in W(Sp)\setminus\{w\}$. Since $w$ and $w^{\prime}$ are swing, there are s-partitions $(K,S),(K^{\prime},S^{\prime})$ for $Sp$ such that
$w\in S,\ w^{\prime}\in S^{\prime},\ |S|=\alpha(Sp)=|S^{\prime}|,\ N(w)=K,\text{ and }\ N(w^{\prime})=K^{\prime}.$
Suppose that $K\neq K^{\prime}$. If $x\in K\setminus K^{\prime}$, then $x\in S^{\prime}$. Hence, $x\in W(Sp)$, and so $K^{\prime}=N(x)=K\setminus\{x\}$. Now, observe that $S^{\prime}=S\dot{\cup}\{x\}$: this is impossible because $S$ and $S^{\prime}$ are maximum independent sets. Therefore, $K=K^{\prime}$, which means that $w^{\prime}$ and $w$ are twins. We have just proven the inclusion $W(Sp)\subset[w]$.
To prove the reverse inclusion, consider a vertex $w^{\prime}$ with $N(w^{\prime})\setminus\{w\}=N(w)\setminus\{w^{\prime}\}$. Since $w$ is swing, there is an s-partition $(K,S)$ for $Sp$ such that $w\in S$ and $N(w)=K$. Then, $N(w^{\prime})=K\setminus\{w^{\prime}\}$, which means that $w^{\prime}\in W(Sp)$. Hence, $w]\subset W(Sp)$. ∎
@@
Since $x\in\operatorname{\mathcal{N}}(G)$, we have
$G_{v*}x=x_{u}+\sum_{i\in N(v)\setminus\{u\}}x_{i}=0=G_{u*}x=x_{v}+\sum_{i\in N(u)\setminus\{v\}}x_{i}.$
Since $N(v)\setminus\{u\}=N(u)\setminus\{v\}$, we have
$\sum_{i\in N(v)\setminus\{u\}}x_{i}=\sum_{i\in N(u)\setminus\{v\}}x_{i}.$
@@
Let $Sp(K,S)$ be an unbalanced split graph. Then, we have the following:
@@
    $W(Sp)$ is a clique or an independent set of mutually twin vertices;
@@
    if $W(Sp)=\{w\}$, then $\operatorname{Supp}(Sp)\subset S^{*}(Sp)\cup w$;
@@
    if $W(Sp)$ is a clique of size $\geq 2$, then $\operatorname{Supp}(Sp)\subset S^{*}(Sp)$; in particular, $W(Sp)\cap\operatorname{Supp}(Sp)=\varnothing$;
@@
    $W(Sp)\subset\operatorname{Supp}(Sp)\subset S^{*}(Sp)\cup W(Sp).$
@@
    It follows from Theorem 4.4, Theorem 4.2(2) and Theorem 4.1(1).